THE FAR-INFRARED RING-PUCKERING ROTATIONAL-VIBRATIONAL SPECTRUM OF 2,5-DIHYDROFURAN

\underline{EUGENE J. BONDOC}, JAAN LAANE, Department of Chemistry, Texas A&M University, College Station, TX 77843.

The far-infrared spectrum of 2,5-Dihydrofuran has been recorded and it shows several ring-puckering transitions along with rotational-vibrational fine structure. Since the molecule is a near symmetric top (A = 0.267 cm^{−1} and B = 0.261 cm^{−1}), the peaks are separated by approximately 0.53 cm^{−1}. Despite the overlap between several puckering bands, the fine structure can be nicely resolved with a resolution of 0.1 cm^{−1}. The ring-puckering spectrum consisting of a series of bands between 99 and 185 cm^{−1} is similar to that previously reported.